# SynthLens: Visual Analytics for Facilitating Multi-step Synthetic Route Design

### v1

@@ I Introduction
Currently, researchers typically use retrosynthesis [8], an automated method that breaks down the target molecule into starting molecules, to construct synthetic routes. However, they often find these routes impractical and prefer to rely on reference papers. Therefore, they explore synthetic reactions from reference papers to design synthetic routes transforming starting molecules into intermediates step by step until obtaining the target molecule. At each step of the designing process, researchers must balance multiple criteria (e.g., yield, duration, and experimental difficulty), which may conflict when selecting the optimal synthetic reaction. Such multi-criteria decision-making task makes the synthetic route design process tedious and time-consuming. Additionally, each decision can affect not only current step’s outcome but also subsequent choices and the overall synthetic route. For instance, choosing the former between two possible reactions at the early step: a high-yield but slower versus a low-yield but fast reaction may improve yield but extend the overall duration, leading to increased costs and delays. The decision may also restrict reagent options in later steps due to intermediate stability. Additionally, the reaction at this early step may limit the choice of reagents in subsequent steps due to the stability of the intermediates produced. The complexity is further heightened by the inclusion of qualitative factors (e.g., experimental procedures) in addition to quantifiable numerical values.
@@ II Problem Formulation
Figure 1: A synthetic route from the starting molecule to the target molecule consists of several synthetic reactions, each involving specific reactants and products. The evaluation of a synthetic reaction contains criteria such as duration and experimental procedure, while.
@@ V-C Information Extraction from Papers
Additionally, we provide users with a key metric: Context Relevancy[42] to help them evaluate the extracted information. Context Relevancy Context Relevancy is an indicator used to evaluate the performance of LLMs in retrieving information, and it measures the relevancy between the content of the answer and its context. Also, to avoid over-reliance on our system, we added disclaimers in SynthLens to remind users that there may be inaccuracy in the extracted information because of technological limitations. Meanwhile, we provide a link to each paper, allowing them to access papers online directly to verify extracted synthetic details.
@@ VI-A2 Paper Projection View
The retrieved papers present a diverse distribution in research content and include detailed information like titles, abstracts, citation counts, keywords, etc., making the exploration of a large number of papers a challenging task. After interviewing experts, we learned they usually rely on heuristic exploration methods to identify target papers for further exploration. This approach allows users to precisely determine the most relevant papers based on a deep understanding of their contents. Therefore, we employ a scatter plot in the Paper Projection View (Fig.3(B)), which helps users find semantically similar papers through the scatter plot after identifying one paper of interest based on their expertise. For example, users are more likely to explore papers spatially close to the selected paper in the scatter plot, rather than just focusing on those ranking highest in the paper list.
@@
After transforming the abstract of these papers into embeddings and dimensionality reduction introduced in Section V, we obtained the coordinates of each paper projected in the 2D plane. However, if these points are mapped directly by coordinates, there will be an overlap among these points. To reduce the overlapping of the points in the scatter plot, we employed a force-directed algorithm provided by d3.js[41] that applies a repulsive force to maintain at least a minimum distance between points, without changing their relative positions. In addition, we added an arc surrounding each point encoding the citation count of this paper, which can indicate its credibility. Proximity in distance among points indicates a similarity in research content. Users are willing to prioritize papers that not only demonstrate a high degree of relevance to the key molecule but also have high citation numbers. Upon exploring a paper, subsequent exploring choices should focus on papers that are near to ensure the consistency of exploration. We display 100 papers in the Paper Projection View because experts usually retrieve less than 100 relevant papers when searching synthetic reactions. Showing too many can overwhelm the analysis with unrelated papers. If the user finds the points too dense, they can zoom in and drag to see a clearer picture of the scatter’s positional relationships.
Interaction. Hovering over a point triggers a tooltip to appear, presenting the details of the corresponding paper, including title, abstract, DOI, citation number, and keywords. These details can guide the users’ decision on whether to explore a specific paper. When a point is clicked, SynthLens will automatically extract the experimental procedures of the expected synthetic reactions from the paper. In addition, we provide a zoom-in button that, when clicked, opens a dialog containing an enlarged scatter plot. In this dialog, we provide two silders: one to adjust the count of displayed points and another to modify the “perplexity” parameter in the t-SNE algorithm, which controls whether the algorithm emphasizes global patterns or local structures. We also include a tooltip to inform users about this parameter.
@@ VI-B3 Rank View
Initially, we presented the scores for three attributes using bar charts. However, based on expert feedback, they felt that presenting multiple bar charts together could lead to visual overload. Therefore, considering the need for ranking, visual simplicity, and quick task completion, we opt for dot plots. Moreover, the human eye is not as sensitive in visual perception to assessing length, i.e., the length of bars, as it is to dots. We use bars with color intensity in the last column to represent the weighted total score. Because the bar chart uses multiple encodings, including the length of the bars, color intensity, and position on the common scale, it can represent the weighted total score while also allowing users to assess the rankings of each row intuitively.
@@ IX Discussion
Design Implications. In our interviews, experts indicated that they can adopt the proposed workflow for designing synthetic routes in practice, treating the process as constructing a tree structure where a node represents a synthetic reaction, and the path from root to leaf defines the synthesis route, with attributes like cumulative yield and synthesis time to help identify the optimal route. Moreover, we observed that experts frequently review intermediate nodes to explore promising synthetic route completions. To facilitate this, SynthLens includes the Molecule Similarity View, allowing users to select any node and automatically receive alternative possibilities of structure-similar molecules. Additionally, users’ feedback indicates that our workflow is suitable for retrosynthetic tasks, where chemists deconstruct a target molecule to find feasible synthetic routes[3]. Our system may also apply to various multi-criteria decision-making scenarios, such as fisheries management[47], architecture construction[48], and clinical decision-making[49]. For instance, in highway route selection, involving multiple objectives like safety, cost etc., our tree-form visualization can help decision-makers compare and rank construction segment plans[50].
@@
Limitations and Future Work. There are several areas that can be addressed to enhance SynthLens further. Firstly, our system supports experts in specifying one currently starting molecule. However, expanding the capability to support multiple starting molecules would give experts more flexibility. Additionally, we can also recommend relevant papers about molecules similar to the current key molecule. This feature is valuable because molecules with similar structures, even if they differ slightly (such as in side chains), may share similar reaction conditions or types. By including these related molecules, the system can help users explore a wider range of relevant synthesis routes. Finally, incorporating additional information from papers, such as spectrograms and validation methods, would enrich the decision-making context for experts. Our main contribution is the proposed analytics analytic pipeline, which allows integration of advanced tools in place of Eunomia, making it both adaptable and scalable to accommodate future advancements.

### v2

@@ I Introduction
Currently, researchers typically use retrosynthesis [8], an automated method that breaks down the target molecule into starting molecules, to construct synthetic routes. However, they often find these routes impractical and prefer to rely on reference papers. Therefore, they explore synthetic reactions from reference papers to manually design synthetic routes transforming starting molecules into intermediates step by step until obtaining the target molecule. At each step of the designing process, researchers must balance multiple criteria (e.g., yield, duration, and experimental difficulty), which may conflict when selecting the optimal synthetic reaction. Such multi-criteria decision-making task makes the synthetic route design process tedious and time-consuming. Additionally, each decision can affect not only current step’s outcome but also subsequent choices and the overall synthetic route. For instance, choosing the former between two possible reactions at the early step: a high-yield but slower versus a low-yield but fast reaction may improve yield but extend the overall duration, leading to increased costs and delays. The decision may also restrict reagent options in later steps due to intermediate stability. Additionally, the reaction at this early step may limit the choice of reagents in subsequent steps due to the stability of the intermediates produced. The complexity is further heightened by the inclusion of qualitative factors (e.g., experimental procedures) in addition to quantifiable numerical values.
@@ II Problem Formulation
Figure 1: A synthetic route from the starting molecule to the target molecule consists of several synthetic reactions, each involving specific reactants and products. The evaluation of a synthetic reaction contains criteria such as duration and experimental procedure.
@@ V-C Information Extraction from Papers
Additionally, we provide users with a key metric: Context Relevancy[42] to help them evaluate the extracted information. Context Relevancy is an indicator used to evaluate the performance of LLMs in retrieving information, and it measures the relevancy between the content of the answer and its context. Also, to avoid over-reliance on our system, we added disclaimers in SynthLens to remind users that there may be inaccuracy in the extracted information because of technological limitations. Meanwhile, we provide a link to each paper, allowing them to access papers online directly to verify extracted synthetic details.
@@ VI-A2 Paper Projection View
The retrieved papers present a diverse distribution in research content and include detailed information like titles, abstracts, citation counts, keywords, etc., making the exploration of a large number of papers a challenging task. After interviewing experts, we learned they usually rely on heuristic exploration methods to identify target papers for further exploration. This approach allows users to precisely determine the most relevant papers based on a deep understanding of their contents. Therefore, we employ a scatter plot in the Paper Projection View (Fig.4(B)), which helps users find semantically similar papers after identifying one paper of interest based on their expertise. For example, users are more likely to explore papers spatially close to the selected paper in the scatter plot, rather than just focusing on those ranking highest in the paper list.
@@
After transforming the abstract of these papers into embeddings and dimensionality reduction introduced in Section V, we obtained the coordinates of each paper projected in the 2D plane. However, if these points are mapped directly by coordinates, there will be an overlap among these points. To reduce the overlapping of the points in the scatter plot, we employed a force-directed algorithm provided by d3.js[41] that applies a repulsive force to maintain at least a minimum distance between points, without changing their relative positions. In addition, we added an arc surrounding each point encoding the citation count of this paper, which can indicate its credibility. Proximity in distance among points indicates a similarity in research content. Users are willing to prioritize papers that not only demonstrate a high degree of relevance to the key molecule but also have high citation numbers. Upon exploring a paper, subsequent exploring choices should focus on papers that are near to ensure the consistency of exploration. We display the 100 papers retrieved as related to the specified molecule in the Paper Projection View because experts usually retrieve less than 100 relevant papers when searching synthetic reactions. Showing too many can overwhelm the analysis with unrelated papers. If the user finds the points too dense, they can zoom in and drag to see a clearer picture of the scatter’s positional relationships.
Interaction. Hovering over a point triggers a tooltip to appear, presenting the details of the corresponding paper, including title, abstract, DOI, citation number, and keywords. These details can guide the users’ decision on whether to explore a specific paper. When a point is clicked, SynthLens will automatically extract the experimental procedures of the expected synthetic reactions from the paper. In addition, we provide a zoom-in button that, when clicked, opens a dialog containing an enlarged scatter plot. In this dialog, we provide two sliders: one to adjust the count of displayed points and another to modify the “perplexity” parameter in the t-SNE algorithm, which controls whether the algorithm emphasizes global patterns or local structures. We also include a tooltip to inform users about this parameter.
@@ VI-B3 Rank View
Initially, we presented the scores for three attributes using bar charts. However, expert feedback indicated that presenting multiple bar charts together could lead to visual overload. Therefore, considering the need for ranking, visual simplicity, and quick task completion, we opt for dot plots. Moreover, the human eye is not as sensitive in visual perception to assessing length, i.e., the length of bars, as it is to dots. We use bars with color intensity in the last column to represent the weighted total score. Because the bar chart uses multiple encodings, including the length of the bars, color intensity, and position on the common scale, it can represent the weighted total score while also allowing users to assess the rankings of each row intuitively.
@@ IX Discussion
Design Implications. In our interviews, experts indicated that they can adopt the proposed workflow for designing synthetic routes in practice, treating the process as constructing a tree structure where a node represents a synthetic reaction, and the path from root to leaf defines the synthesis route, with attributes like cumulative yield and synthesis time to help identify the optimal route. Moreover, we observed that experts frequently review intermediate nodes to explore promising synthetic route completions. To facilitate this, SynthLens includes the Molecule Similarity View, allowing users to select any node and automatically receive alternative possibilities of structure-similar molecules. Additionally, users’ feedback indicates that our workflow is suitable for retrosynthetic tasks, where chemists deconstruct a target molecule to find feasible synthetic routes[3]. Our method may also apply to various multi-criteria decision-making scenarios, involving multiple options at each stage and enables users to efficiently evaluate the impact of choices on the outcome, such as fisheries management[47], architecture construction[48], and clinical decision-making[49]. For instance, in highway route selection, involving multiple objectives like safety, cost etc., our tree-form visualization can help decision-makers compare and rank construction segment plans[50].
@@
Limitations and Future Work. There are several areas that can be addressed to enhance SynthLens further. Firstly, our system supports experts in specifying one currently starting molecule. However, expanding the capability to support multiple starting molecules would give experts more flexibility. Additionally, we can also recommend relevant papers about molecules similar to the current key molecule. This feature is valuable because molecules with similar structures, even if they differ slightly (such as in side chains), may share similar reaction conditions or types. By including these related molecules, the system can help users explore a wider range of relevant synthesis routes. Finally, incorporating additional information from papers, such as spectrograms and validation methods, would enrich the decision-making context for experts. Our main contribution is the proposed analytics analytic pipeline, which allows integration of advanced tools in place of our method, making it both adaptable and scalable to accommodate future advancements.